OpenStudy (marigirl):

simplify 4th root of 36

OpenStudy (marigirl):

\[\sqrt[4]{36}\]

ganeshie8 (ganeshie8):

As a start change it to exponent form using : \[\large \sqrt[n]{m}= m^{\frac{1}{n}}\]

OpenStudy (marigirl):

3*3*2*2

OpenStudy (marigirl):

how do i end up at sqrt6 as my answer

OpenStudy (zzr0ck3r):

\((\sqrt{6}\sqrt{6})(\sqrt{6}\sqrt{6})=6*6\) Go with @ganeshie8 's idea to see how you can get there without knowing..

OpenStudy (marigirl):

\[(36)^{1/4}\] \[6^2(^.25)\]

OpenStudy (marigirl):

then we can go 6 ^ 0.5

OpenStudy (marigirl):

which is square root of 6

OpenStudy (marigirl):

thanks @ganeshie8 @zzr0ck3r

OpenStudy (zzr0ck3r):

Yeah, but what is the strategy if we do not know what we are looking for ? @ganeshie8

OpenStudy (marigirl):

i worked it out on my calc and found that it was square root 6. for some reason i was looking for one number that can be multiplied by 4 times to get to 36 which is not the case in scenario